OpenStudy (anonymous):

giving medal for answer !!!!!, HELP Find the distance between the two points. Round to the nearest tenth if necessary. (10, –8), (4, 0) a. 2 b. 10 c. 16 d. 100

OpenStudy (anonymous):

distance formula is \[\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (anonymous):

woah...idk how to do all that

OpenStudy (anonymous):

when you plug the numbers in it will look like this \[\sqrt{(4-10)^{2}+(0-(-8)^{2}}\] first subtract 4-10 and 0-(-8) the square them answers then add them together then take the square root

OpenStudy (anonymous):

4-10=-6->6^2=36 0-(-8)=8->8^2=64 64+36=100 the square root of 100 is 10